nomencl^{∗}

A Package to Create a Nomenclature

Bernd Schandl^{†}

                          schandl@gmx.net

Printed on January 31, 2000

1 Introduction

How often did you try to understand a theorem in a book, but just couldn’t figure out what all those strange symbols were all about? The nomencl package should help authors format a nomenclature. It uses the powerful capabilities of the MakeIndex program to generate such a list automatically using information provided by the author throughout the text.

2 Package Options

The nomencl package has the following options:

refeq The phrase “, see equation (⟨eq⟩)” is appended to every entry in the nomenclature where ⟨eq⟩ is the number of the last equation in front of the corresponding command \nomenclature.

norefeq No equation reference is printed. (default)

refpage The phrase “, page ⟨page⟩” is appended to every entry in the nomenclature where ⟨page⟩ is the number of the page on which the corresponding command \nomenclature appeared.

norefpage No page reference is printed. (default)

prefix Every sort key is preceded by the letter “a” (changeable); see Section 4 to learn why this might make sense. (default)

noprefix No prefix is used. danish, english, french, german, italian, polish, russian, spanish,

∗Package version v2.9 of 1999/11/23.

†Up to version v2.2 this package was maintained by Boris Veytsman.

ukrainian The reference texts and the nomenclature title will appear in the corresponding language. Note that in order to use Russian or Ukrainian, you have to have Cyrillic fonts installed and you might need a replacement for MakeIndex, e. g. xı^{∘}ndy. Please help me out with other languages. (default: english)

3 Usage

3.1 The Basics

The creation of the nomenclature list is very similar to the creation of an index [6, App. A]. You need to:

• Put \usepackage[⟨options⟩]{nomencl} in the preamble of your document.

\makeglossary • Put \makeglossary in the preamble of your document.

\nomenclature • Issue the \nomenclature command (see Section 3.2) for each symbol you want to have included in the nomenclature list. The best place for this command is immediately after you introduce the symbol for the first time.

\printglossary • Put \printglossary at the place you want to have your nomenclature list.

Now put your file through L^{A}TEX. The command \makeglossary will instruct L^{A}TEX to open the glossary file ⟨filename⟩.glo corresponding to your L^{A}TEX file ⟨filename⟩.tex and to write the information from your \nomenclature commands to this file.

The next step is to invoke MakeIndex. You should instruct MakeIndex to use ⟨filename⟩.glo as your input file, use nomencl.ist as your style file^{1}, and write output to the file ⟨filename⟩.gls. How to do this depends on your implementation of MakeIndex. For most UNIX implementations you should write something like

   makeindex ⟨filename⟩.glo -s nomencl.ist -o ⟨filename⟩.gls

Now you have the file ⟨filename⟩.gls that contains your nomenclature list properly ordered. The last step is to invoke L^{A}TEX on your master file ⟨filename⟩.tex once more. It will input your .gls file and process it accordingly to the current options. That’s all!

1German users who want to use the shortcut notation "a instead of \"a have to redefine the quote character in nomencl.ist to something other than " (and |, @, !), maybe &; see the comment in the source code section and in the file nomencl.ist.

3.2 The Main Command

\nomenclature The main command of the nomencl package has the following syntax:

\nomenclature[⟨prefix⟩]{⟨symbol⟩}{⟨description⟩} where ⟨prefix⟩ is used for fine tuning the sort order, ⟨symbol⟩ is the symbol you want to describe and ⟨description⟩ is the actual description. The sortkey will be ⟨prefix⟩⟨symbol⟩, where ⟨prefix⟩ is either the one from the optional argument or, if no optional argument was given, the default ⟨prefix⟩ which may be empty. See Section 4 to make sense of this.

Put this command immediately after the equation or text that introduces ⟨symbol⟩. Usually it is a good idea to avoid a space or an unquoted newline just in front of the \nomenclature command. Put a % at the end of the preceding line if necessary. Don’t forget to enclose math in ⟨symbol⟩ in $ signs.

Let’s have a look at a simple example. If your input file looks like the one in Figure 1 then your nomenclature^{2} should look like Figure 2.

 \documentclass{article}
 \usepackage{nomencl}
 \makeglossary
 \begin{document}
 \section*{Main equations}
 \begin{equation}
   a=\frac{N}{A}
 \end{equation}%
 \nomenclature{$a$}{The number of angels per unit area}%
 \nomenclature{$N$}{The number of angels per needle point}%
 \nomenclature{$A$}{The area of the needle point}%
 The equation $\sigma = m a$%
 \nomenclature{$\sigma$}{The total mass of angels per unit area}%
 \nomenclature{$m$}{The mass of one angel}
 follows easily.
 \printglossary
 \end{document}

Figure 1: Input of a simple example

Note the necessary quoting of newlines. When the \nomenclature macros appear directly after the equation environment, quote all newlines; when they appear in the middle of a sentence, quote all but the last newlines.^{3}

2Note that all the examples are somewhat faked in this document, but they give a good impression of the “real” result.

3I’m not sure how to resolve this more elegantly, but suggestions are welcome.

Nomenclature

σ The total mass of angels per unit area

A The area of the needle point

a The number of angels per unit area

m The mass of one angel

N The number of angels per needle point

Figure 2: Output of a simple example

Due to the way \nomenclature scans its arguments you don’t need to \protect any macros, but you also must not have any character in front of the first or between the first and the second argument, especially no line break (even with a %). So

   \nomenclature{$x$}%
     {Description}

does not work. You can have have line breaks in the argument, but also no %.

3.3 Referencing

\refeq As explained in Section 2, you can turn referencing to equations and pages

\refpage
\refeqpage
\norefeq
\norefpage
\norefeqpage

on/off globally using the package options. But sometimes you might want to change the referencing behavior for single entries. The following six macros can be used inside a \nomenclature macro: \refeq, \norefeq,

\refpage, \norefpage, \refeqpage, \norefeqpage. The first four work

similar to the package options, only local to the entry; the last two are shortcuts, but saying \refeqpage is equivalent to \refeq\refpage.

If we change the relevant parts of the last example as shown in Figure 3 then the nomenclature should look like Figure 4.

While these macros do not have to be at the end of the entries, it’s probably the most sensible place to put them. Note that such local request always supersede the package options.

\begin{equation}
  a=\frac{N}{A}
\end{equation}%
\nomenclature{$a$}{The number of angels per unit area\refeqpage}%
\nomenclature{$N$}{The number of angels per needle point\refeq}%
\nomenclature{$A$}{The area of the needle point\refeq\refpage}%
The equation $\sigma = m a$%
\nomenclature{$\sigma$}{The total mass of angels per unit area}%
\nomenclature{$m$}{The mass of one angel\refpage}
follows easily.
\printglossary
\end{document}

Figure 3: Input with references

Nomenclature

σ The total mass of angels per unit area

A The area of the needle point, see equation (1), page 1

a The number of angels per unit area, see equation (1), page 1

m The mass of one angel, page 1

N The number of angels per needle point, see equation (1)

Figure 4: Output with references

4 Sort Order of the Entries

The Greek letter σ turned out to be first in the nomenclature list in the examples above because the backslash in \sigma precedes any alphabetical character. Sometimes this is not what you want. Then you can use ⟨prefix⟩ to fine tune the sort order.

Before we describe the usage of ⟨prefix⟩, we have to explain how MakeIndex sorts entries, see [2]. MakeIndex distinguishes three kinds of sort keys:

Strings Everything that starts with a alphabetic letter (A… Z, a… z).

Numbers Everything that starts and only contains digits (0… 9).

Symbols Everything else.

Each group is sorted separately (and differently), then the groups are sorted in the order symbols, numbers, strings. For the groups the following algorithm^{4} is used:

Strings If two letters are compared, the usual ordering is used (a<C<q), but if two words are the same except for the capitalization, then an upper case letter precedes the lower case letter (Tea<tea). If a letter is compared with a non-letter (digit, symbol), ASCII code is used (1<A, A<~).^{5} If two non-letters are compared (which can not happen at the first position of a string), ASCII code is used (+<1<:<\). Additionally there is the issue of word ordering (treat spaces as letter with ASCII code smaller than every printable symbol) and letter ordering (ignore spaces). MakeIndex uses word ordering by default, but you can change it with some command line option (-l on my UNIX).

Numbers The natural ordering is used (8<34<111).

Symbols ASCII code is used (+<1<:<A<\<a).

Why did you have to read all this?^{6} Let’s consider the following eight nomenclature entries (without the optional argument): $~Ab$, $~aa$, $\Ab$, $\aa$, $Ab$, $aa$, Ab, aa. Try to understand the following example with the help of the explanation above and an ASCII table.

If you use nomencl with its default settings (i. e. “a” is added to every sort key, so every sort key is considered as a string), you will get the sort

4This is only vaguely described in [2], so I had to figure out special cases by myself.

Please correct me if I am wrong

5An exception seems to be that the non-letters between upper and lower case letters (code 91–96) are put just before the capital letters (between code 64 and 65) while the non-letters after the lower case letters (code 123–127) are left there. Can someone please enlighten me why?

6I hope you did read it ;-)

order $\aa$, $\Ab$, $aa$, $Ab$, $~aa$, $~Ab$, aa, Ab. Note that aa is in

front of Ab in all four pairs; note also the order $\Ab$, $Ab$, $~Ab$ which does not agree with the ASCII code.

If you specify the option noprefix, then you will get $Ab$, $\Ab$, $\aa$, $aa$, $~Ab$, $~aa$, aa, Ab. The first six entries are considered as symbols and sorted according to the ASCII code (this time correctly). Note that $\Ab$ is in front of $\aa$ because A has the smaller ASCII code. The two strings follow at the end.

Decide for yourself what you prefer. Personally, I like to specify the noprefix option and use the optional argument to get exactly the sort order I want. See Section 6 for some special effects.

5 Customization

Besides the things you can customize by using the package options there are a few more commands that you might want to redefine. If you make the same changes in every file, it’s probably easier to put all those in a file nomencl.cfg which is automatically read by the nomencl package whenever it exists in the search path.

5.1 Formatting the Nomenclature

\printglossary Probably the most common change to the nomenclature is a different \nomlabelwidth amount of space for the symbols. By default, the nomenclature is for-

matted as a list with the label width equal to \nomlabelwidth which is initialized to 1 cm. You can change this dimension in the cfg file or you can use the optional argument of \printglossary. If you want to have a little more space for the labels (and you don’t live in a metric world) you can use

   \printglossary[0.5in]

instead of the simple

   \printglossary

theglossary If you don’t like the format of the nomenclature at all, you will have to redefine the theglossary environment. Maybe a look at the documented code of nomencl will help.

\nomname In case you don’t like the name of the nomenclature, just redefine the

\nomname macro, e. g.

   \renewcommand{\nomname}{List of Symbols}

If you are using e. g. the documentclass book with page style headings you should also take care of correct headings:

   \cleardoublepage% or \clearpage
   \markboth{\nomname}{\nomname}% maybe with \MakeUppercase
   \printglossary

I thought about putting this in the definition of \printglossary but decided that it is much easier for the user to add it if he wants than to remove it if he doesn’t want it. In case you always need this just define a macro in nomencl.cfg that executes these three lines all at once and can be used

instead of \printglossary.

\nomgroup Usually, MakeIndex inserts the macro \indexspace between every character group, i. e. between symbols and numbers, numbers and letters and between every two letter groups. The nomencl package inserts the macro \nomgroup{⟨arg⟩} instead, where ⟨arg⟩ is either the string “Symbols” or the string “Numbers” or the capital letter of the group that is about to start. You can redefine \nomgroup to insert some white space

   \renewcommand{\nomgroup}[1]{\medskip}

or to print a fancy divider

   \renewcommand{\nomgroup}[1]{%
     \item[]\hspace*{-\leftmargin}%
     \rule[2pt]{0.45\linewidth}{1pt}%
     \hfill #1\hfill
     \rule[2pt]{0.45\linewidth}{1pt}}

Note that \nomgroup is executed in a list environment, so you need to have an \item first and then jump back to the beginning of the line with the \hspace command.

\nompreamble Maybe you want to explain something just between the title of the \nompostamble nomenclature and the start of the list or at the very end of the list. Just

redefine the macros \nompreamble and \nompostamble which do nothing by default. Note that they are executed outside of the list environment.

\nomitemsep The skip between two entries in the nomenclature can be adjusted using \nomitemsep. This should be done in the preamble or the file nomencl.cfg. Note that if you want no extra skip between entries you have to use

   \setlength{\nomitemsep}{-\parsep}

\nomprefix If you want, you can redefine the default prefix that is used for the sortkeys. By default, \nomprefix is set to “a”; redefining it supersedes the package options prefix and noprefix.

5.2 Formatting the Entries

\nomlabel By default, the labels are just shifted to the left within their allocated box. If you want to change this, redefine \nomlabel which should get one argument, e. g.

   \renewcommand{\nomlabel}[1]{\hfil #1\hfil}

to center the symbols.

\nomentryend Maybe you would like to have a period at the end of every entry. Just

say

   \renewcommand{\nomentryend}{.}

and there it is. Section 6.2 explains another nice application of this macro.

\eqdeclaration If you don’t like the text that is used for the references to equations

\pagedeclaration and pages, you can define \eqdeclaration and \pagedeclaration. Both

should accept one argument, namely the equation and page number, respectively. An example is

   \renewcommand{\eqdeclaration}[1]{, first used in eq.~(#1)}

If you are redefining these macros for a particular language, let me know and I will add that language to the next release of the nomencl package.

6 Tips and Tricks

In this section, I will gather fancy stuff that people did or might want to do with the nomencl package. Please email any ideas you have.

6.1 Subgroups

If you have distinct groups among the identifiers in your nomenclature (e. g. Greek letters for physical constants, Roman letters for variables) you can use the optional argument of \nomenclature together with the \nomgroup macro to get two groups with separate headings in the nomenclature.

Use something like the following throughout your text

   \nomenclature[ga ]{$\alpha$}{Constant}
   \nomenclature[rx ]{$x$}{Variable}

where “g” and “r” indicate Greek and Roman letters, respectively. Then you include the ifthen package and redefine \nomgroup e. g. like this:

    \renewcommand{\nomgroup}[1]{%
      \ifthenelse{\equal{#1}{R}}{\item[\textbf{Variables}]}{%
        \ifthenelse{\equal{#1}{G}}{\item[\textbf{Constants}]}{}}{}}

Note that we have to check for capital letters. All your symbols should have some kind of prefix; maybe you can also use the default prefix “a”. Note that for symbols and numbers you have to check for the strings “Symbols” and “Numbers”.

6.2 Units

Besides the obvious possibility of adding units for symbols in the description string, you can also use \nomentryend to shift the unit to the right margin. Something along the lines of a macro

    \newcommand{\nomunit}[1]{%
      \renewcommand{\nomentryend}{\hspace*{\fill}#1}}

should do the job. You can use this macro like this

   \nomenclature{$l$}{Length\nomunit{m}}

Note that the nomenclature will not be a tabular with three columns, but it is pretty close as long as you only have one-line descriptions. Any suggestions for improvements are welcome.

6.3 Using a Long Table instead of a List

The following idea was sent to me by Brian Elmegaard. I have modified it a little bit to make it work with the current version of nomencl. Only the basic idea is given, so you have to do some extra thinking (and coding) to get it to work the way you want it.

After loading the longtable package in the preamble we first have to modify the macro that writes the entries to the glo file (do this is an style file).

    \def\@@@nomenclature[#1]#2#3{%
     \def\@tempa{#2}\def\@tempb{#3}%
     \protected@write\@glossaryfile{}%
      {\string\glossaryentry{#1\nom@verb\@tempa @{\nom@verb\@tempa}&%
          \begingroup\nom@verb\@tempb\protect\nomeqref{\theequation}%
            |nompageref}{\thepage}}%
     \endgroup
     \@esphack}

Then the nomenclature itself must be changed to start a longtable instead of a list. Maybe we could add something for a repeating header on every page.

    \def\theglossary{%
      \@ifundefined{chapter}{\section*}{\chapter*}{\nomname}%
      \nompreamble
      \begin{longtable}[l]{@{}ll@{}}}
    \def\endtheglossary{%
      \end{longtable}%
      \nompostamble}

Finally we add the following two lines at the end of nomencl.ist^{7}.

    item_0  ""
    delim_t " \\\\\n"

As I said, this is only the basic idea. An advantage might be the repeating headers on every page, a disadvantage is that there won’t be any line breaks in the second column.

7Don’t forget to rename the file and delete my email address if you want to distributed the file, see the pointer to the LPPL in Section 8.

7 Acknowledgements

First and foremost I want to thank Boris Veytsman, who had the idea for the package, maintained it until v2.2 and provided some helpful advice for the new version. I also want to thank Stefan Böhm and Karl Heinz Marbaise who helped testing this package.

For helping out with translations I thank Brian Elmegaard (Danish), Artur Gorka (Polish), Denis B. Roegel (French), Alejandro Lopez-Valencia (Spanish), Boris Veytsman (Russian and Ukrainian) and Sani Egisto (Italian).

8 Releases and Legal Issues

This package can be redistributed and/or modified under the terms of the L^{A}TEX Project Public License distributed from CTAN archives in the directory macros/latex/base/lppl.txt, see e. g. [3]; either version 1 of the license, or (at your option) any later version.

The most recent release of the nomencl package can always be found at

http://www.math.clemson.edu/~bschand/nomencl/. Usually, the same
version is also available at CTAN/macros/latex/contrib/supported/
nomencl/.

References

[1] Braams, Johannes; Carlisle, David; Jeffrey, Alan; Lamport, Leslie; Mittelbach, Frank; Rowley, Chris; Schöpf, Rainer (1996). ltidxglo.dtx – 1996/01/20 v1.1e LaTeX Kernel (Index and Glossary). CTAN/macros/

   latex/base/ltidxglo.dtx.

[2] Chen, Pehong; Harrison, Michael A. (1987). Automating Index Preparation. Report UCB/CSD 87/347, Computer Science Division, University of California, Berkeley, CA.

[3] Comprehensive TEX Archive Network CTAN. ftp://ctan.tug.org/

   tex-archive/.

[4] Jones, David M. (1995). A new implementation of L^{A}TEX’s indexing commands, Version v4.1beta of 1995/09/28. CTAN/macros/latex/

   contrib/supported/camel/index.dtx.

[5] Knuth, Donald E. (1984). The TEXbook. Addison-Wesley Publishing Company, Reading, MA.

[6] Lamport, Leslie (1994). L^{A}TEX: A Document Preparation System. Addison-Wesley Publishing Company, Reading, MA.

[7] Veytsman, Boris (1996). Package nomencl, Version 2.2. http://www.

   plmsc.psu.edu/~boris/nomencl/.